OpenStudy (agl202):

Medal For help... -> Solve for x, given the equation Square root of x minus 5+ 7 = 11. x = 21, solution is extraneous x = 21, solution is not extraneous x = 81, solution is extraneous x = 81, solution is not extraneous

OpenStudy (agl202):

@ganeshie8 hlp.. plz...

Nnesha (nnesha):

\[\huge\rm \sqrt{x-5}+7=11\] like this ?

Nnesha (nnesha):

or 7 is also under the root ?

OpenStudy (agl202):

So it would be x=21 right?

OpenStudy (agl202):

So, A.?

Nnesha (nnesha):

what is an extraneous and non-extraneous solution means ?

OpenStudy (agl202):

Extraneous Solution is a solution of the simplified form of an equation that does not satisfy the original equation.

Nnesha (nnesha):

okay. substitute x for its value into the original equation if both sides are equal then 21 isn't extraneous solution if both sides aren't equal then 21 is an extraneous solution

Nnesha (nnesha):

\[\huge\rm \sqrt{x-5}+7=11\] replace x with 21 solve left side

OpenStudy (agl202):

I don't get it. ^

Nnesha (nnesha):

x=21 right now to check if it's an extraneous or non-extraneous you should substitute x for 21

OpenStudy (agl202):

I get a non-extraneous.

Nnesha (nnesha):

right \[\sqrt{21-5}+7=11\] both sides are equal so it would be non-ex

OpenStudy (agl202):

You deserve a medal! Thx u so much!!! :D

Nnesha (nnesha):

yw!

OpenStudy (agl202):

I got it right!!

Nnesha (nnesha):

nice